OpenStudy (anonymous):

You have already invested $550 in a stock with an annual return of 11%. How much of an additional $1,400 should be invested at 12% and how much at 6% so that the total return on the entire $1,950 is 9%? (Round each answer to the nearest cent.)

OpenStudy (helder_edwin):

Let x the amount invested at 12% and y at 6% then solve \[ \large \begin{cases} x+y=1400\\ .12x+.06y+550\times.11=1950\times.09 \end{cases} \]

OpenStudy (anonymous):

Do you substitute?

OpenStudy (helder_edwin):

solve it the way u feel more comfortable with.

OpenStudy (helder_edwin):

though the easiest would be substitution.

OpenStudy (anonymous):

Ok thank you

OpenStudy (helder_edwin):

u r welcome